Grade 12th passMechanics

13 days is the half-life period of a sample. After how many days, the sample will become 1/16th of the original substance

Abin Santhosh
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
k = ln2 / t1/2
now find k and then use 
k = (1/t) ln(a/ at )
at = a/6

Regards
Arun (askIITians forum expert)